OpenStudy (anonymous):

MEDAL What are the solutions of 3x2 + 4x = -5? the quantity negative 4 plus or minus i square root of 66 all over 6 the quantity negative 2 plus or minus 2i square root of 66 all over 3 the quantity negative 2 plus or minus i square root of 11 all over 3 the quantity negative 4 plus or minus i square root of 11 all over 6

OpenStudy (anonymous):

@khusker

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

Have you studied the quadratic equation in your class?

OpenStudy (danjs):

\[x = \frac{ (-b) +- \sqrt{b^2 - 4 * a * c} }{ 2a }\] where \[ax^2 + bx + c = 0\]

OpenStudy (danjs):

Your equation is: 3x^2 + 4x + 5 = 0, so a = 3 b = 4 c = 5

OpenStudy (danjs):

using those values, in the quadratic formula above ...

OpenStudy (anonymous):

i'm confused lol

OpenStudy (anonymous):

ohhhhh

OpenStudy (danjs):

\[x = \frac{ -4 +- \sqrt{4^2 - 4*3*5} }{ 2*3 }\]

OpenStudy (danjs):

The square root is going to contain a negative number, meaning imaginary/complex numbers

OpenStudy (anonymous):

i got -44

OpenStudy (danjs):

right, recall the square root of -1 is i

OpenStudy (danjs):

\[\sqrt{-44} = \sqrt{-1 * 4 * 11} = \sqrt{-1}\sqrt{4}\sqrt{11} \]

OpenStudy (danjs):

\[2i * \sqrt{11}\]

OpenStudy (anonymous):

so would that be -2 square root of 11

OpenStudy (danjs):

2i square root 11

OpenStudy (anonymous):

i though i equals -1. or is that i squared

OpenStudy (anonymous):

*thought

OpenStudy (danjs):

\[\sqrt{-1} = i\]

OpenStudy (danjs):

so \[i^2 = -1\]

OpenStudy (anonymous):

soo i'm guessing it's C?

OpenStudy (danjs):

\[x = \frac{ -4 +- 2i \sqrt{11} }{ 6 } = \frac{ -2 +- i \sqrt{11} }{ 3 }\]

OpenStudy (anonymous):

yes!! i was right lol

OpenStudy (danjs):

just divided each term by 2 there to reduce it

OpenStudy (anonymous):

ohkayy. thank you so much

OpenStudy (danjs):

so it is the 3rd answer

OpenStudy (danjs):

no prob, anytime